OpenStudy (anonymous):

please explain: 5x^2-30x=0

OpenStudy (anonymous):

i'm guessing you want to solve for x. factor out the common term(s). in this case: 5x \[5x^2 - 30x = 0\]\[5x(x - 6) = 0\] for this to hold true, by the zero-product rule, either 5x = 0 or (x - 6) = 0 therefore the answer is x = 0 OR x = 6

OpenStudy (anonymous):

did you do a simplified FOIL when you did 5x(x-6)=0

OpenStudy (anonymous):

and what step did you do after that to get 6?

OpenStudy (anonymous):

i guess you can look at it that way. i factored out 5x since they both have 5x in common. and i looked at it as 5x = 0 therefore x = 0 and x - 6 = 0 , added 6 to both sides, x = 6

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

glad i could help :)